### v1

@@ Abstract
The reawakened era of lunar exploration is defined by a strategic shift from temporary visits to a sustained international and commercial presence, resulting in an unprecedented demand for a robust and continuously available communication infrastructure. The conventional direct-to-Earth communication architecture relies on limited and oversubscribed deep space networks, which are further challenged by the radiative environment and insufficient visibility in certain areas of the cislunar domain. We address these issues by proposing a foundational move toward inter-domain space network cooperation by introducing architectures based on near space networks. They can directly service lunar surface users or, via cislunar relays, by forming a resilient and multi-layered communication backbone. First, we establish a unified link analysis framework incorporating frequently disregarded environmental factors, such as the Moon’s variable illumination, to provide a high-fidelity performance evaluation. Second, we assess architectures’ reliability based on the outage risk, essential for quantifying the operational robustness of communication links. Finally, to manage the inherent dynamism of architectures, we propose an inter-domain space digital twin—an dynamic decision-making engine that performs real-time analysis to autonomously select the best communication path, ensuring high and stable reliability while simultaneously optimizing power consumption. Overall, our paper provides a holistic architectural and conceptual management framework, emphasizing the necessity of lunar communications to support a permanent human and economic foothold on the Moon.
@@ 1 INTRODUCTION
The 21st century marks a new era of lunar exploration by shifting from temporary visits to establishing a permanent and sustainable presence on the Moon [1]. A diverse ecosystem of commercial companies and international partners affiliating with long-term ambitions for settlement and resource utilization are all part of this revitalized global endeavor [2]. This strategic shift from transient missions to sustained presence necessitates continuous operations, a multitude of interacting assets [3], including habitats and rovers [4], and sophisticated logistical chains [5]. Therefore, a continuous and reliable communication backbone connecting the Moon to Earth without any interruptions is a vital requirement for these missions, as it has a direct impact on each phase of the mission [6].
@@
The conventional way of lunar communication is to establish a direct-to-Earth (DTE) link, which is advantageous due to its simplicity and lower latency [7]. Despite its simplicity, this solution is accompanied by limitations and challenges that cannot be overlooked, making the conventional communication architecture mismatched to the demands of the new lunar era. The recent attempts by Japan’s iSpace highlight these factors that led to the loss of contact with the landers and their eventual crash on the Moon’s surface [8], despite both missions being close to achieving a soft landing. Particularly, the visibility of certain areas on the Moon from a ground station significantly varies during a day, resulting in limited line-of-sight (LoS) availability [9]. To mitigate this, NASA operates three deep space network (DSN) facilities strategically located about apart, but even with this global distribution, any single location on the near side of the Moon has limited and intermittent contact windows [10]. This problem is more critical for the lunar far side, which is permanently shielded from Earth, making DTE communication unfeasible without a relay [11].
@@
Establishing a DSN facility is also not economically scalable. To overcome the immense path loss due to the distance, ground stations must be equipped with advanced systems, such as giant antennas [12] (such as those with diameters of 70 meters) and high-power amplifiers that transmit signals in the kilowatt range [13]. This implies that constructing and operating a DSN station can incur costs in the hundreds of millions of dollars, which is why increasing the number of DSN stations for direct communication is not an easily applicable option. Additionally, the DSN is a heavily oversubscribed resource and stations are not always used for lunar missions, as they are essentially intended to support deep space missions, creating regulatory constraints [14] and scheduling conflicts [15]. Therefore, access to DTE links is challenging for non-agency users such as commercial companies or academic institutions. As these challenges become more apparent with the increasing number of missions, cislunar space networks (CSNs) have been planned by both commercial and national actors to reduce costs, expand access, and ensure reliable communication [16].
@@
This new paradigm shifts away from centrally managed point to point links to a flexible [17], scalable [18], and interoperable [19] network of networks such as NASA’s emerging framework, LunaNet [20]. These solutions are designed to provide the robust and on-demand connectivity required to sustain a permanent presence on the lunar surface by burgeoning multi-layered constellations of relay satellites [21] and advanced networking protocols [22]. CSN architectures in orbits such as low lunar orbit (LLO) [23] or near rectilinear halo orbit are proposed to overcome connectivity problems at the Moon [24]. However, while CSNs ensure local coverage even for assets on the far side or in shadowed polar regions, a critical architectural question remains: how do these networks create a robust and efficient transmission back to Earth? This stimulates inter-domain space networks by facilitating state-of-the-art cooperation among multi-layered CSNs and near space networks (NSNs).
@@
To advance this cooperation between CSNs and NSNs, a more versatile but rational analysis framework is required, given that both cislunar and near space domains111Near space domain is the operational region extending from the Karman line [25] to geostationary orbit (GEO) at km while the cislunar space domain is the volume from GEO to approximately km beyond the Moon’s second Lagrange point [3]. have different operational considerations, such as intense traffic in near space [26] or quite limited sources in cislunar space [27]. The high congestion of NSNs, especially at lower orbits, and related problems [28] make the concept of a GEO network with two-way communication links more reasonable for inter-domain space networks, as they also provide more stable and wide coverage in both directions. However, the need for multifaceted and generalized link analysis models still exists, while current models are often insufficient because their oversimplified structure fails to capture the unique aspects of each domain and the complex interactions between them [2]. A critical oversight in many models is the impact of variable thermal noise [29], where the changing brightness temperature of the lunar surface directly contributes to the system noise temperature of a receiver [30] and impacts link reliability by degrading the signal-to-noise ratio (SNR), especially at higher frequency bands like Ka-band [31, 32, 33]. This emphasizes the need of a unified link analysis framework for inter-domain space communication, which embodies these refined environmental effects to provide an accurate and holistic architectural assessment.
@@
The holistic assessments to quantify and mitigate the risk of communication failure under dynamic real-time conditions are mandatory [34], as a link failure for high stakes operations, especially those involving crewed missions or expensive robotic assets, is not an acceptable option. This situation elevates the importance of rigorous reliability assessment and motivates the outage probability as a critical performance metric to analyze the operational robustness and trustworthiness of critical systems [35]. Furthermore, the performance of any given communication architecture—whether DTE, LLO relay, or a hybrid LLO-GEO system—is not always optimal due to the dynamic interplay of orbital mechanics, network topology, and environmental conditions throughout the missions. Therefore, relying solely on architectural designs for inter-domain space networks is not entirely satisfactory, which raises a compelling need for dynamic management of these architectures. Such a system must dynamically select the best communication path—switching between lunar communication architectures—to prioritize link reliability while also accounting for domain-level operational limitations like power consumption [36]. We therefore propose an inter-domain space digital twin. Its decision-making engine uses a high-fidelity virtual model of the network, updated with real-time telemetry, to predict performance, diagnose faults, and autonomously command the network to adapt its routing configuration, ensuring the most reliable and efficient communication path is always active.
@@
We addresses the severe limitations of conventional lunar communication, which hinder continuous access and mission scalability, by proposing a comprehensive collaboration of inter-domain space networks. We introduce two architectures that integrate a GEO network (in the near space domain) with a LLO network (in the cislunar space domain), enabling more robust and cost-effective communication essential for future lunar endeavors.
@@
Overcoming inherent trade-offs in lunar communication architectures and ensuring seamless, high performance communication, we introduce an inter-domain space digital twin. This decision-making engine moves beyond static network designs by offering a dynamic optimization approach that continuously assesses real-time link conditions. It dynamically selects the best scenario by prioritizing reliability while also optimizing power utilization, which is critical for managing the complexity of emerging lunar architectures and enabling autonomous and robust missions.
We validate our contributions with presented numerical results that are obtained with the simulations in a well-known digital mission engineering tool, Satellite Tool Kit (STK) [37]. The analysis of in-orbit dynamics across four communication scenarios reveals significant fluctuations in link geometry and visibility, highlighting the inherent unreliability of each architecture. Reliability assessments, based on outage probability, quantify the performance of each architecture, demonstrating that while introduced NSN integrated architectures usually offer advantages, no single option provides continuous, robust connectivity under all dynamic conditions. The proposed inter-domain space digital twin successfully maintains a stable and highly reliable communication link, even under challenging and variable noise environments, by dynamically selecting the best scenario based on real-time link conditions and expected reliability targets.
@@ 2 UNIFIED FRAMEWORK FOR LINK ANALYSIS
To assess the suggested lunar communication architectures, we need a detailed analysis system that can represent the complicated and changing environment at different space domains. Conventional link models are frequently insufficient, as their simple structure fails to represent the distinctive features of the cislunar and near space realms. This section, therefore, develops a unified link analysis framework that holistically integrates critical link budget factors to offer accurate analysis. This comprehensive approach provides the essential foundation for the precise performance and reliability assessments conducted in the subsequent sections.
@@ 2.1.1 Path Loss
The free space path loss is given as
@@ 2.1.3 Atmospheric Loss
The atmospheric loss () is calculated by considering the gaseous attenuation in accordance with the recommendations by the international telecommunication union as below [38],
@@ 2.2 Thermal Noise
Thermal noise arises from the thermal emissions of electronic devices. It is superimposed on the transmitted signal at the receiver and typically modeled with additive white Gaussian noise (AWGN). To determine SNR, we must characterize the thermal noise sources, which are collectively represented by an operational equivalent noise temperature . The total noise power is given as follows [29, 30]
@@ 2.2.2 Antenna Noise Temperature
The external antenna noise temperature, is aroused from the brightness temperature of the subtended body and calculated [39] as follows
Here, is the antenna beam solid angle [40] as
where . denotes the normalized antenna power pattern. In this paper, it is defined for the antennas with the diameter [41] as below
@@
The reason for this correction is that the noise sources in radio astronomy exhibit random polarization, which causes the antenna to receive only half of the power [40]. The solid angle at the distance subtended by the bright body (Moon or Earth) is for any in steradian as follows
where denotes the radius of the bright body that is km and km for Moon and Earth, respectively. , the internal antenna noise temperature depends on the physical temperature of the antenna and is calculated as below
@@ 2.3 Signal and Channel Models
The calculation of the received signal and noise power requires the consideration of different factors through different architectural scenarios depending on the source and destination: Moon (M), LLO, GEO, Earth (E). For the scenario , the received signal between the source and the destination is defined as follows
@@
All links between a transmitter on the lunar surface to a receiver at any location are subject to Rician fading [42] and the cumulative distribution function of the SNR is as follows:
@@ 3 RELIABILITY ASSESSMENT FOR LUNAR COMMUNICATION ARCHITECTURES
Building upon the link analysis framework, this section quantifies the reliability of lunar communication architectures for inter-domain space network cooperation. We utilize the outage probability for each network architecture as the reliability performance metric since failures or substantial deterioration in communication quality might compromise mission objectives and even threaten assets or human lives. For any given communication link of -th scenario is given depending on the minimum SNR threshold () and instantaneous SNR () as [43]
@@
For the second scenario, communication occurs from the Moon to a LLO relay, and then from the LLO relay to Earth ground station. As a failure of either the LLO-to-Earth link or the Moon-to-LLO link results in an outage for the scenario, the outage probability is expressed as:
@@
The fourth scenario involves a multi-hop path: Moon-to-LLO, then LLO-to-GEO, and finally GEO-to-Earth. The outage probability is computed as:

### v2

@@ Abstract
The resurgent era of lunar exploration is defined by a strategic shift from temporary visits to a sustained international and commercial presence, resulting in an unprecedented demand for a robust and continuously available communication infrastructure. The conventional direct-to-Earth communication architecture relies on limited and oversubscribed deep space networks, which are further challenged by the radiative environment and insufficient visibility in certain areas of the cislunar domain. We address these issues by proposing a foundational move toward inter-domain space network cooperation by introducing architectures based on near space networks. They can directly service lunar surface users or, via cislunar relays, by forming a resilient and multi-layered communication backbone. First, we establish a unified link analysis framework incorporating frequently disregarded environmental factors, such as the Moon’s variable illumination, to provide a high-fidelity performance evaluation. Second, we assess architectures’ reliability based on the outage risk, essential for quantifying the operational robustness of communication links. Finally, to manage the inherent dynamism of architectures, we propose an inter-domain space digital twin: a dynamic decision-making engine that performs real-time analysis to autonomously select the best communication path, ensuring high and stable reliability while simultaneously optimizing power consumption. Overall, our paper provides a holistic architectural and conceptual management framework, emphasizing the necessity of lunar communications to support a permanent human and economic foothold on the Moon.
@@ 1 INTRODUCTION
The 21st century marks a new era of lunar exploration by shifting from temporary visits to establishing a permanent and sustainable presence on the Moon [CH7_1]. A diverse ecosystem of commercial companies and international partners affiliated with long-term ambitions for settlement and resource utilization are all part of this revitalized global endeavor [Selen1]. This strategic shift from transient missions to sustained presence necessitates continuous operations, a multitude of interacting assets [CH2_5], including habitats and rovers [ref5_t1], and sophisticated logistical chains [2ndsub_1]. Therefore, a continuous and reliable communication backbone connecting the Moon to Earth without any interruptions is a vital requirement for these missions, as it has a direct impact on each phase of the mission [CH7_2].
@@
The conventional way of lunar communication is to establish a direct-to-Earth (DTE) link, which is advantageous due to network simplicity and lower latency [ref1_t1]. Despite its simplicity, this solution is accompanied by limitations and challenges that cannot be overlooked, making the conventional communication architecture mismatched to the demands of the new lunar era. The recent attempts by Japan’s iSpace highlight these factors that led to the loss of contact with the landers and their eventual crash on the Moon’s surface [Japan], despite both missions being close to achieving a soft landing. Particularly, the visibility of certain areas on the Moon from a ground station significantly varies during a day, resulting in limited line-of-sight (LoS) availability [CH7_3]. To mitigate this, NASA operates three deep space network (DSN) facilities strategically located about apart, but even with this global distribution, any single location on the near side of the Moon has limited and intermittent contact windows [DSN]. This problem is more critical for the lunar far side, which is permanently shielded from Earth, making DTE communication unfeasible without a relay [baris1].
@@
Establishing a DSN facility is also not economically scalable. To overcome the immense path loss due to the distance, ground stations must be equipped with advanced systems, such as giant antennas [ref4_t1] (such as those with diameters of 70 meters) and high-power amplifiers that transmit signals in the kilowatt range [R1]. This implies that constructing and operating a DSN station can incur costs in the hundreds of millions of dollars, which is why increasing the number of DSN stations for direct communication is not an easily applicable option. Additionally, the DSN is a heavily oversubscribed resource, and the stations are not always used for lunar missions, as they are essentially intended to support deep space missions, creating regulatory constraints [CH7_4] and scheduling conflicts [CH7_5]. Therefore, access to DTE links is challenging for non-agency users such as commercial companies or academic institutions. As these challenges become more apparent with the increasing number of missions, cislunar space networks (CSNs) have been planned by both commercial and national actors to reduce costs, expand access, and ensure reliable communication [Selen2].
@@
This new paradigm shifts away from centrally managed point-to-point links to a flexible [CH7_6], scalable [CH7_7], and interoperable [ref2_t1] network of networks such as NASA’s emerging framework, LunaNet [LunaNet]. These solutions are designed to provide the robust and on-demand connectivity required to sustain a permanent presence on the lunar surface by burgeoning multi-layered constellations of relay satellites [CH7_8] and advanced networking protocols [CH7_9]. CSN architectures in orbits such as low lunar orbit (LLO) [baris2] or near rectilinear halo orbit are proposed to overcome connectivity problems at the Moon [CH7_10]. However, while CSNs ensure local coverage even for assets on the far side or in shadowed polar regions, a critical architectural question remains: how do these networks create a robust and efficient transmission back to Earth? This stimulates inter-domain space networks by facilitating state-of-the-art cooperation among multi-layered CSNs and near space networks (NSNs).
@@
To advance this cooperation between CSNs and NSNs, a more versatile but rational analysis framework is required, given that both cislunar and near space domains111Near space domain is the operational region extending from the Karman line [CH2_1] to geostationary orbit (GEO) at km while the cislunar space domain is the volume from GEO to approximately km beyond the Moon’s second Lagrange point [CH2_5]. have different operational considerations, such as intense traffic in near space [CH2_2] or quite limited sources in cislunar space [CH7_11]. The high congestion of NSNs, especially at lower orbits, and related problems [C1R30] make the concept of a GEO network with two-way communication links more reasonable for inter-domain space networks, as they also provide more stable and wide coverage in both directions. However, the need for multifaceted and generalized link analysis models still exists, while current models are often insufficient because their oversimplified structure fails to capture the unique aspects of each domain and the complex interactions between them [Selen1]. For example, while space links are designed for a dominant LoS, the Moon’s craterous topography introduces significant multipath effects from scattering and reflection, creating a complex channel that must be accurately characterized for reliable link design. Additionally, the harsh radiative environment of deep space, influenced by factors like solar wind and plasma, can degrade signal quality and is often overlooked. A critical oversight in many models is the impact of variable thermal noise [kt1], where the changing brightness temperature of the lunar surface directly contributes to the system noise temperature of a receiver [kt2] and impacts link reliability by degrading the signal-to-noise ratio (SNR), especially at higher frequency bands like Ka-band [C5R26, ref7_t1, ref8_t1]. This emphasizes the need for a unified link analysis framework that, unlike prior art which often relies on conventional link budgets or models isolated phenomena, holistically integrates these refined environmental effects to provide an accurate and holistic architectural assessment.
@@
The holistic assessments to quantify and mitigate the risk of communication failure under dynamic real-time conditions are mandatory [C4R3], as a link failure for high-stakes operations is not an acceptable option. This urgency was dramatically highlighted during Astrobotic’s Peregrine mission last year; although a critical propulsion anomaly prevented a lunar landing, its robust communication link was the sole enabler for the ground team to diagnose the problem, salvage valuable data from onboard payloads, and ultimately guide the spacecraft to a safe, controlled re-entry [Mission_fail2]. This situation elevates the importance of rigorous reliability assessment and motivates the outage probability as a critical performance metric to analyze the operational robustness and trustworthiness of critical systems [sennur]. Furthermore, the performance of any given communication architecture—whether DTE, LLO relay, or a hybrid LLO-GEO system—is not always optimal due to the dynamic interplay of orbital mechanics, network topology, and environmental conditions throughout the missions. Therefore, relying solely on static architectural designs for inter-domain space networks is not entirely satisfactory, which raises a compelling need for dynamic management of these architectures.
While dynamic management is a mature concept in terrestrial networks, its application to the unique multi-domain (CSN-NSN) cislunar environment, with its stringent reliability demands and complex, high-latency link variations, represents a significant, unaddressed challenge. Such a system must dynamically select the best communication path—switching between lunar communication architectures—to prioritize link reliability while also accounting for domain-level operational limitations like power consumption [CH7_12]. We therefore propose an inter-domain space digital twin. Its decision-making engine uses a high-fidelity virtual model of the network, updated with real-time telemetry, to predict performance, diagnose faults, and autonomously command the network to adapt its routing configuration, ensuring the most reliable and efficient communication path is always active.
@@
We address the severe limitations of conventional lunar communication, which hinder continuous access and mission scalability, by proposing a comprehensive collaboration of inter-domain space networks. We introduce two architectures that integrate a GEO network (in the near space domain) with a LLO network (in the cislunar space domain), enabling more robust and cost-effective communication essential for future lunar endeavors.
@@
To overcome inherent trade-offs in lunar communication architectures and ensure seamless, high-performance communication, we introduce an inter-domain space digital twin. This decision-making engine moves beyond static network designs by offering a dynamic optimization approach that continuously assesses real-time link conditions. It dynamically selects the best scenario by prioritizing reliability while also optimizing power utilization, which is critical for managing the complexity of emerging lunar architectures and enabling autonomous and robust missions.
We validate our contributions with presented numerical results that are obtained with the simulations in a well-known digital mission engineering tool, Satellite Tool Kit (STK) [STK_Ansys]. The analysis of in-orbit dynamics across four communication scenarios reveals significant fluctuations in link geometry and visibility, highlighting the inherent unreliability of each architecture. Reliability assessments, based on outage probability, quantify the performance of each architecture, demonstrating that while introduced NSN integrated architectures usually offer advantages, no single option provides continuous, robust connectivity under all dynamic conditions. Conversely, we demonstrate that the proposed inter-domain space digital twin’s dynamic orchestration resolves this gap, reducing total daily downtime by over 99.9% compared to the best-performing lunar architecture alternative. It achieves this by successfully maintaining a stable and highly reliable communication link, even under challenging and variable noise environments, through dynamically selecting the best scenario based on real-time link conditions and expected reliability targets.
@@ 2 UNIFIED FRAMEWORK FOR LINK ANALYSIS
To assess the suggested lunar communication architectures depicted in Fig. 1, we need a detailed analysis system that can represent the complicated and changing environment at different space domains. Conventional link models are frequently insufficient, as their simple structure fails to represent the distinctive features of the cislunar and near space realms. This section, therefore, develops a unified link analysis framework that holistically integrates critical link budget factors to offer accurate analysis. This comprehensive approach provides the essential foundation for the precise performance and reliability assessments conducted in the subsequent sections.
@@ 2.1.1 Path Loss
The free space path loss is calculated for each one-way communication link in the end-to-end path as follows
@@ 2.1.3 Atmospheric Loss
The atmospheric loss () is calculated by considering the gaseous attenuation in accordance with the recommendations by the international telecommunication union as below [ITU_P676_13],
@@ 2.2 Thermal Noise
Thermal noise arises from the thermal emissions of electronic devices. It is superimposed on the transmitted signal at the receiver and typically modeled with additive white Gaussian noise (AWGN). To determine SNR, we must characterize the thermal noise sources, which are collectively represented by an operational equivalent noise temperature . The total noise power is given as follows [kt1, kt2]
@@ 2.2.2 Antenna Noise Temperature
The external antenna noise temperature, arises from the brightness temperature of the subtended body and calculated [R3_ITU] as follows
Here, is the antenna beam solid angle [R2_kraus] as
where . denotes the normalized antenna power pattern. In this paper, it is defined for the antennas with the diameter [STK_antenna] as below
@@
The reason for this correction is that the noise sources in radio astronomy exhibit random polarization, which causes the antenna to receive only half of the power [R2_kraus]. The solid angle at the distance subtended by the bright body (Moon or Earth) is for any in steradian as follows
where denotes the radius of the bright body that is km and km for the Moon and Earth, respectively. , the internal antenna noise temperature depends on the physical temperature of the antenna and is calculated as below
@@ 2.3 Signal and Channel Models
The calculation of the received signal and noise power requires the consideration of different factors through different architectural scenarios depending on the source and destination: Moon (M), LLO, GEO, Earth (E). For the architectural scenario , the received signal between the source and the destination is defined as follows
@@
All links originating from the lunar surface are modeled using the Rician fading distribution, as it accurately captures the channel characteristics: a strong, dominant LoS path coexisting with multipath components generated by signal scattering off the Moon’s craterous terrain [C5R34]. The Rician -factor provides a physically intuitive parameter to represent the ratio of power between the direct LoS component and these scattered components. While other fading models exist, the Rician distribution offers a more accurate representation than the Rayleigh model, which assumes no LoS path [Multipath1]. It also provides a better balance of accuracy and tractability [RiceK4] for this LoS-dominant scenario [multipath2] compared to more generalized models like the distribution, a choice well-supported by prior work in the field [Multipath1, multipath2, RiceK4, RiceK1]. The cumulative distribution function of the SNR for a Rician channel is given as:
@@ 3 RELIABILITY ASSESSMENT FOR LUNAR COMMUNICATION ARCHITECTURES
Building upon the link analysis framework, this section quantifies the reliability of lunar communication architectures in Fig. 1 for inter-domain space network cooperation. We utilize the outage probability for each network architecture as the reliability performance metric since failures or substantial deterioration in communication quality might compromise mission objectives and even threaten assets or human lives. For any given communication link of -th architecture’s scenario, the outage probability is given depending on the minimum SNR threshold () and instantaneous SNR () as [alouini]
@@
In relaying scenarios, the satellites function as active transponders that utilize the decode-and-forward technique [DF1]. Therefore, the signal is fully regenerated at each hop, making the communication links statistically independent. The total outage probability is thus calculated as the probability that at least one of these independent links fails, causing an outage for the end-to-end chain. For the second scenario, communication occurs from the Moon to a LLO relay, and then from the LLO relay to Earth ground station. As a failure of either the LLO-to-Earth link or the Moon-to-LLO link results in an outage for the scenario, the outage probability is expressed as:
@@
The fourth scenario involves a multi-hop path: Moon-to-LLO, then LLO-to-GEO, and finally GEO-to-Earth. The end-to-end outage probability is determined by subtracting the joint success probability of all three links from one as: